OpenStudy (anonymous):

Explain how you would find the exact value of the square root of 2 over the square root of 8.

OpenStudy (anonymous):

\[\sqrt2/\sqrt8 = \sqrt{2/8} \]by the quotient rule.

OpenStudy (anonymous):

I know but im thinking it is 1 over 4 cause if you put 2/8 into simplest for you will get 1/4

OpenStudy (anonymous):

Close, but it still needs to be square rooted (that sounds weird). Luckily, 1/4th can be square rooted easily.

OpenStudy (anonymous):

@ julucero so you multiply bot by 2 right

OpenStudy (anonymous):

Why would you multiply the bottom by 2?

OpenStudy (anonymous):

which is 1/2 @amishjeb

OpenStudy (anonymous):

Bingo. :)

OpenStudy (anonymous):

lol :D

OpenStudy (anonymous):

So.....2 sqrt over sqrt 8 equals 1/2 because you multiply 1/4 by 2 to simplify it?!?!?

OpenStudy (anonymous):

Also\[\frac{ \sqrt{2} }{ \sqrt{8} } = \frac{ \sqrt{2} }{ 2\sqrt{2} } = \frac{ 1 }{ 2 } \times \frac{ \sqrt{2} }{ \sqrt{2} } = \frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

ohh I was close

OpenStudy (anonymous):

It happens to get you the right answer in this case, but its not right. You need to square 1/4th.\[\sqrt{1} = 1 and \sqrt{4} = 2\]

OpenStudy (anonymous):

ohhh okay I get it if we use 1/4 it would of been wrong I think

OpenStudy (anonymous):

well you can do this another way technically square root means 2 _____ / x ^2 = x^2

OpenStudy (anonymous):

Well. Sort of. Maybe? I'm having a hard time understanding what you're asking. But taking the square root isn't the same as multiplying by 2.

OpenStudy (anonymous):

yea I know well thanks anyways for the help

OpenStudy (anonymous):

but you still wind up getting the same answer :/

OpenStudy (anonymous):

i said the same thing but it is whatever @mth3v4

OpenStudy (anonymous):

after all that typing i dont think i kept track lol

OpenStudy (anonymous):

lol :D